OpenStudy (anonymous):

In 1990, people in the United States spent about $285.7 billion on recreation. In 1995, they spent $402.5 billion. Using a linear equation in slope-intercept form to model the amount (in billions of dollars) spent on recreation since 1990, predict the amount of money spent on recreation in the United States in 2000.

OpenStudy (amistre64):

we can make this simpler by -1990 and making that 0,285.7

OpenStudy (amistre64):

(5,402.5) -(0,285.7) ---------- (5 ,116.8) ; slope = y/x = 116.8/5

OpenStudy (anonymous):

OP posted it twice

OpenStudy (anonymous):

look back 3 posts from this one, and you get the same question

OpenStudy (amistre64):

y = 116.8/5(x-1990) when x = 2000; y = 233.6 right?

OpenStudy (amistre64):

yep, its there :)

OpenStudy (anonymous):

YOUR AWSOME (: thanks so much (:

OpenStudy (amistre64):

yw :)